# Automated Review Generation Method Based on Large Language Models

### v1

@@ Abstract
Literature research, vital for scientific advancement, is overwhelmed by the vast ocean of available information. Addressing this, we propose an automated review generation method based on Large Language Models (LLMs) to streamline literature processing and reduce cognitive load. In case study on propane dehydrogenation (PDH) catalysts, our method swiftly generated comprehensive reviews from 343 articles, averaging seconds per article per LLM account. Extended analysis of 1041 articles provided deep insights into catalysts’ composition, structure, and performance. Recognizing LLMs’ hallucinations, we employed a multi-layered quality control strategy, ensuring our method’s reliability and effective hallucination mitigation. Expert verification confirms the accuracy and citation integrity of generated reviews, demonstrating LLM hallucination risks reduced to below 0.5% with over 95% confidence. Released Windows application enables one-click review generation, aiding researchers in tracking advancements and recommending literature. This approach showcases LLMs’ role in enhancing scientific research productivity and sets the stage for further exploration.
@@ 1 Introduction
In scientific research, peer-reviewed academic literature serves as a dense and reliable medium for information dissemination, enabling researchers to push the boundaries of human knowledge by building on previous work[1]. The clarity and rigor of scientific language constrain information dissemination, making it a key carrier in the research process for entity description, concept extraction, information transfer, and consensus building. This ensures that in the transmission and evolution of knowledge, both the sender and receiver of information construct highly consistent models of the referenced objects and concepts on the cognitive level. For instance, the development of industrial catalysts requires a thorough understanding of the materials’ structure, chemical properties, and reactivity, considering their activity, selectivity, and stability[2, 3, 4, 5, 6]. This necessitates leveraging foundational catalytic theories and reaction mechanisms detailed in literature. However, the rapid pace of literature publication has outstripped researchers’ capacity to assimilate information[7, 8], highlighting the need for tools to efficiently analyze and integrate data, thus avoiding redundant discoveries and broadening research perspectives, ultimately facilitating scientific research processes such as catalyst development.
Natural language processing (NLP), a branch of machine learning (ML), with core tasks like co-reference resolution and semantic analysis[9], is well-suited as such a tool for literature understanding and integration. NLP technology has evolved through feature engineering, neural network architecture engineering, pre-trained language model development, and prompt-based learning[10]. Recent years have seen NLP applied in catalysis literature for extracting synthetic methods[11, 12, 13, 14, 15, 16], materials and properties[11, 12, 13, 14, 17, 18, 19, 20, 21, 22], key reaction parameters[11, 16, 19, 23, 24, 25, 26], and reactions[14, 24, 27, 28]. However, these studies often focus on isolated aspects, which limits their transferability and requires prior domain knowledge and programming skills, making it challenging for newcomers. Naturally, we considered integrating literature information in the form of reviews, which can also be generalized to more disciplines. However, early attempts at automated review generation often treated it as a multi-document summarization (MDS) task[29], relying on existing reviews and citation networks[30, 31, 32, 33, 34], thus may not keep pace with rapid advancements in research. Additionally, bibliometric analyses can overlook recent studies due to insufficient citations; focusing solely on abstracts or citations rather than full texts[31, 32, 33, 34] may miss critical details. Focusing on extractive summarization rather than integrated generation[30, 31, 32], or filling in sentence templates[33] can omit significant information or be redundant.
Since November 2022, Large Language Models (LLMs) like ChatGPT, representing breakthroughs in NLP, have demonstrated unprecedented language comprehension abilities[35]. LLMs excel in zero-shot and few-shot learning, common sense, logical reasoning, and versatility across NLP tasks[36], serving as a second brain for researchers that processes and comprehends extensive scientific literature without additional foundational knowledge[35, 37]. However, LLMs’ creative reprocessing of understood information differs from search engines and poses the challenge of "hallucinations", a prevalent but unresolved issue in the industry[37, 38, 39]. Research reveals that calibrated LLMs inevitably generate hallucinations[40], a phenomenon unavoidable in any computable LLM, regardless of model architecture, learning algorithm, prompt techniques, or training data[41]. Hallucinations in LLMs, which often emerge from statistical biases or noise in the training data and strategies for handling ambiguous information, manifest as baseless false information, contextually misaligned or unrelated responses[38]. Hallucinations are exacerbated in specialized domains by limited data exposure, leading LLMs to produce seemingly credible yet misleading outputs with specialized terminology, posing significant risks in scientific research where accuracy is critical. Even advanced LLM like GPT-4 achieve only 73.3% accuracy in professional contexts[42], risking inaccurate academic conclusions and misdirected research, causing considerable time and resource losses[39]. Ensuring hallucination mitigation is crucial for the scientific integrity and reliability of automated review generation. Galactica[43], which claimed review generation capabilities, was withdrawn due to hallucination concerns, highlights this necessity. Therefore, this study emphasizes mitigating hallucinations and enhancing reliability to safeguard the quality of LLM-generated reviews.
Addressing existing methodological limitations and leveraging LLMs’ potential, this study develops an LLM-based, efficient, and comprehensive automated review generation approach. It features an end-to-end pipeline for literature retrieval, reading, summary distillation, and coherent text organization, underpinned by a multi-tier quality control strategy to counter LLM hallucination risks. LLMs’ adeptness at information refinement and knowledge building enables accurate entity and concept extraction from texts, bolstering literature handling, refinement, and knowledge construction capabilities of researchers. By exploiting computational strengths in storage and parallel processing, LLMs overcome human cognitive constraints, easing researchers’ cognitive load and aiding in the rapid identification of research trends. This automated review generation method thus presents a panoramic field view, offering a one-click solution for researchers lacking relevant background. Below is a comparative analysis of automated and traditional review generation methods (see Table 1):
Table 1: Comparison of automated review generation method and traditional literature review method
Feature Automated Review Generation Method Traditional Literature Review Method
Processing Speed Efficient, completing reading of a literature within seconds per LLM account Inefficient, requiring several hours for each document
Scalability Easily scalable by adding more LLM accounts Difficult to increase speed by adding more experienced personnel
Human Resources Saves time for professionals Requires substantial professional manpower
The automated review generation method based on LLMs holds substantial scientific significance in research. It enhances literature processing efficiency and quality, fosters new knowledge discovery, and stimulates innovation. This method is invaluable in advancing the scope and depth of contemporary scientific research. As an innovative literature processing tool, it could become integral to scientific research infrastructure, significantly promoting scientific research progress.
2 Results
@@ 2.1 Automated retrieval
In our study, automated review generation essentially reprocesses the retrieved information. The process hinges on efficiently retrieving and extracting pertinent information from extensive scientific literature, with the review’s quality and scope directly tied to the retrieval process’s comprehensiveness and accuracy. We utilized SerpAPI for automated retrieval on Google Scholar, focusing on propane dehydrogenation (PDH) catalysts, covering literature from 1980 to 2024 in top-tier(Q1) chemistry and chemical engineering journals according to the 2022 Chinese Academy of Sciences division table.
The automated retrieval yielded 1420 initial results from Google Scholar. To address the challenge of irrelevant or duplicate findings, we implemented a dual-level filtering process. The first level employed quick filtering of abstracts and titles to remove obviously irrelevant documents, serving as a rapid but less precise narrowing method. The second level involved deeper LLM-based analysis of full texts, offering higher accuracy albeit at a slower pace. This coarse-to-fine screening method, reminiscent of high-throughput screening, enabled us to efficiently and accurately identify literature pertinent to our research. The initial screening shortlisted 343 articles as related to our topic. Subsequent LLM evaluation further confirmed 238 of these articles as relevant.
@@ 2.2 Implementation and analysis of one-click automated review generation
Using PDH catalysts as an example and building on the aforementioned automated retrieval, we have effectively produced high-quality, specialized review articles. By focusing on top-tier journals, we ensured the retrieval of articles with significant academic impact, offering an accessible starting point for users new to the domain. For those with domain familiarity, the program allows the specification of a custom journal list to refine article selection.
We evaluated the efficacy of this method by contrasting two strategies for constructing review topics: one based on existing reviews and another using LLM-generated topics (see Table 2). The examples showcased in subsequent sections and the Supplementary Information are based on outlines derived from existing reviews.
Table 2: Comparison of review topic construction methods
Method Number of topics Number of guiding questions Number of citations Prior knowledge required
Review-based topic generation 9 35 125 Yes
LLM automated topic generation 12 12 43 No
Key benefits of the generated reviews, irrespective of the topic construction method, include:
  1. 1.
    Content Accuracy: The content has been manually checked by experts in the relevant field, with no errors in knowledge, correct referencing of cited literature, and a length and citation count that align with conventional review standards. Specific examples of generated content can be found in the Supplementary Information.
  2. 2.
    Customizable Research: Enables the addition of specific questions to tailor research focus and refine review specificity.
  3. 3.
    Forward-Looking Insights: Each topic includes a section on "Comprehensive understanding and prospective outlook", providing profound insights and innovative suggestions by LLM. This allows the LLM to engage in divergent thinking and integration beyond the information domain provided by the literature.
To improve accessibility, we developed a Python3 graphical user interface (GUI), enabling straightforward, one-click review generation on Windows, requiring no programming skills or domain knowledge.
@@ 2.3 Data mining and visual analysis
In this study, the data mining module was deployed for comprehensive analysis in the PDH catalysts domain, examining literature from 1980 to 2024 within the chemistry and chemical engineering journals ranked Q1, Q2, and Q3 by the 2022 Chinese Academy of Sciences. Out of 1041 articles filtered by abstracts and titles, 839 were pinpointed as pertinent to PDH research via LLM selection. Leveraging LLMs for data extraction and subsequent analysis, we provided insightful conclusions on catalysts’ composition, structure, and performance. This approach not only highlighted PDH research trends but also explored the maximum performance of individual factors and the synergistic effects between variables.
For instance, a statistical analysis of the annual publication numbers by catalyst types (see Figure 1 (a)) and sources of performance enhancement (see Figure 1 (b)) showed a surge in alloy research since 1995 and a spike in single-atom catalyst studies post-2015, and primarily driven by advancements in structural composition. This trend underscores the PDH field’s evolving focus and hints at fresh avenues for catalyst development, including synthesis methods. In our analysis of the impact of promoter elements (see Figure 1 (c)) and support materials (see Figure 1 (d)) on catalyst performance, including selectivity and stability, we identified that promoter elements such as Zn, Sn, and La, as well as support materials like alumina and zeolites, can achieve notable peak performance, which signaled pathways for catalytic innovation. The combination analysis, for instance, of active site elements with composition elements (see Figure 1 (e)) and alloy structure types with preparation methods (see Figure 1 (f)), revealed that multi-metal systems generally outperform single-metal systems, especially when promoter elements like Sn, Zn, In are used to enhance the performance of Pt-based catalysts. Moreover, impregnation-prepared nanometallic catalysts exhibited superior conversion rates and selectivity, while single-atom alloys showed high selectivity but lower conversion rates.
This comprehensive analysis reveals the nuanced interplay between variables, guiding future research towards optimizing catalyst performance, aiding researchers in achieving the optimal performance balance in catalyst design and optimization. It suggests selecting Pt-based catalysts for maximum selectivity or metal oxides for enhanced conversion rates, and conducting deeper exploration into single-atom and nanostructured catalysts, which show promise in exceeding the efficacy of conventional catalysts. These insights not only showcase the diverse characteristics and performance benchmarks within the PDH domain but also highlight LLMs’ utility in scientific exploration, providing researchers with real-time domain understanding and progress perception, thereby fostering catalyst development. This holistic approach empowers researchers to refine catalyst design and optimization effectively, aligning with industrial needs.
[…]

### v3

@@ Abstract
Literature research, vital for scientific work, faces the challenge of surging information volumes exceeding researchers’ processing capabilities. We present an automated review generation method based on large language models (LLMs) to overcome efficiency bottlenecks and reduce cognitive load. Our statistically validated evaluation framework demonstrates that the generated reviews match or exceed manual quality, offering broad applicability across research fields without requiring users’ domain knowledge. Applied to propane dehydrogenation (PDH) catalysts, our method swiftly analyzed 343 articles, averaging seconds per article per LLM account, producing comprehensive reviews spanning 35 topics, with extended analysis of 1041 articles providing insights into catalysts’ properties. Through multi-layered quality control, we effectively mitigated LLMs’ hallucinations, with expert verification confirming accuracy and citation integrity while demonstrating hallucination risks reduced to below 0.5% with 95% confidence. Released Windows application enables one-click review generation, enhancing research productivity and literature recommendation efficiency while setting the stage for broader scientific explorations.
Keywords: large language models, automated review generation, literature analysis, scientific writing
@@ 1 INTRODUCTION
Peer-reviewed academic literature functions as a critical medium for scientific knowledge dissemination, enabling researchers to advance human understanding through cumulative progress[1]. The clarity and rigor of scientific language facilitates entity description, concept extraction, and consensus building, ensuring cognitive consistency between information senders and receivers. However, the exponential growth in publications has exceeded researcher’ processing capacity[2, 3, 4], necessitating efficient tools for literature analysis and integration, thus avoiding redundant discoveries and broadening research perspectives.
Natural language processing (NLP), encompassing co-reference resolution, semantic analysis, etc.[5], powers literature comprehension. Since November 2022, Large Language Models like ChatGPT, as the latest NLP advancement, have exhibited unprecedented language understanding[6]. Leading LLMs have surpassed human performance on various benchmarks including MMLU[7], which tests undergraduate knowledge, and GPQA Diamond[8], which assesses graduate-level reasoning, positioning them as potential "second brains" for efficient literature processing[6, 9]. Recent studies like PaperQA[10] and its improved version PaperQA2[11], as retrieval-augmented generation (RAG[12]) agents, demonstrate excellent performance in literature-related tasks including retrieval, question-answering, summarization, and contradiction detection, surpassing human expertise in some aspects. AcademicGPT[13] provides comprehensive research support, while CuriousLLM[14] enhances multi-document question-answering through reasoning-based traversal agents. However, these applications require user-provided literature, rely on question-answer interactions, or focus on specific points, limiting their transferability.
The review format effectively integrates literature information and generalizes across disciplinary fields, naturally leading to automated review generation research. However, early attempts encountered several limitations. They either reduced reviews to multi-document summarization[15] or depended on existing reviews and citation networks[16, 17, 18, 19, 20], which struggled with rapidly evolving fields and underestimated recent publications due to citation lag. Additional constraints included using only abstracts instead of full texts as input data[17, 18, 19, 20] and employing either extractive summarization rather than integrated generation[16, 17, 18] or template-based generation[19], risking information loss and redundancy. Recent LLM-based solutions include: multi-AI agent systems[21] for full-process automation from research question generation to data extraction; LitLLM[22] combining RAG with LLM reranking to generate high-quality literature reviews based on user-provided abstracts; LLAssist[23] and related work[24] for literature screening; and ChatCite[25] improving summary quality through human-like workflows. These advances enhance automated review generation while enabling efficient academic research.
Based on the potential of LLMs, this study proposes an automated review generation method based on LLMs, builds an end-to-end data pipeline from literature retrieval to final review text generation. By leveraging information refinement and knowledge construction capabilities of LLMs, this method overcomes human cognitive limitations in single-threaded processing and memory capacity, reducing researchers’ cognitive load while offering superior speed and scalability, thereby substantially conserving professional human resources. However, two critical challenges persist: the macro-level requirement for systematic quality evaluation and comparison with manual reviews, and the micro-level necessity to effectively mitigate LLM hallucinations. To address these challenges, we designed a dual-baseline automatic evaluation framework with rigorous statistical validation, alongside multi-level quality control strategies throughout the process. The distinctive feature of method lies in its adaptability to diverse disciplinary terminologies and knowledge structures without domain-specific training, facilitating both comprehensive field overviews for experienced researchers and accessible entry points for those lacking relevant background, opening up new possibilities for promoting interdisciplinary research and knowledge dissemination. This approach holds substantial scientific significance by enhancing literature processing efficiency, fostering knowledge discovery, and stimulating innovation. Its ability to promote interdisciplinary communication and knowledge integration positions it as a potential cornerstone of modern research infrastructure, accelerating scientific discovery and technological advancement across domains.
2 RESULTS AND DISCUSSION
@@ 2.1 Automated retrieval
Automated review generation fundamentally relies on retrieving and extracting scientific literature, with output quality dependent on input timeliness, quality, and breadth. To demonstrate cross-disciplinary generalization without human intervention, we conducted a case study on propane dehydrogenation (PDH) catalysts, searching chemistry and chemical engineering journals (1980-2024) ranked Q1 in the Chinese Academy of Sciences journal classification on Google Scholar through SerpAPI.
The automated retrieval yielded 1420 initial results from Google Scholar. To address the challenge of irrelevant or duplicate findings, we implemented a dual-level filtering process. The first level employed quick filtering of abstracts and titles to remove obviously irrelevant documents, as detailed in Method section, serving as a rapid but less precise narrowing method. The second level involved deeper LLM-based analysis of full texts, offering higher accuracy albeit at a slower pace. This coarse-to-fine screening method, reminiscent of high-throughput screening, enabled us to identify literature pertinent efficiently and accurately to our research. The initial screening shortlisted 343 articles as related to our topic. Subsequent LLM evaluation further confirmed 238 of these articles as relevant.
@@ 2.2 Implementation and analysis of one-click automated review generation
Using PDH catalysts as an example and building on the aforementioned automated retrieval, we have effectively produced high-quality, specialized review articles. Considering that the entire process is completely end-to-end without the need for human intervention, we believe that a single domain example is sufficient to demonstrate the applicability of this method. The main reason for limiting the journal range to Q1 journals is that although the impact factor of journals may not be closely related to the quality of articles, the lower limit of literature in Q1 journals that have passed strict peer review is relatively higher. Considering users lacking prior knowledge in the target domain, directly traversing Q1 chemistry journals provides an efficient starting point. We solemnly declare here that we are not encouraging users to only consider Q1 journals, but rather suggesting that in the initial stage, one can consider starting with Q1 journals, and for research on the entire field, all possibilities should be explored, which is also supported by our method. In the Windows GUI we provide, using Q1/Q2-3 journals is an optional button, allowing users to choose for themselves. For those with domain familiarity, the program allows the specification of a custom journal list to refine article selection.
We evaluated two topic construction strategies: based on existing reviews (9 topics, 35 questions, 125 citations) and direct LLM generation (12 topics, 12 questions, 43 citations). The examples showcased in subsequent sections and SI are based on outlines derived from existing reviews. The content has been manually checked by experts in the relevant field, with no errors in knowledge, correct referencing of cited literature, and a length and citation count that align with conventional review standards (see SI). The method’s effectiveness stems from LLMs’ human-level or superior language comprehension abilities, coupled with the injection of domain knowledge from retrieved literature through context window, thereby enabling generalization across all research fields. Beyond content accuracy, the method enables customizable research focus through supporting of adding specific questions and provides forward-looking insights with comprehensive understanding sections. To facilitate broad adoption, we developed an open-source Python3 GUI enabling one-click review generation without programming expertise or domain knowledge.
2.3 Evaluation of generated review quality
Research demonstrates LLMs excel in evaluation tasks, with GPT-4 surpassing both crowdsourced workers[26] and experts[27] in text annotation accuracy and reliability, and bias control for complex tasks requiring contextual knowledge reasoning[27]. LLMs show comparable or superior performance to human annotation in persuasiveness, accuracy, and satisfaction[28]. Studies confirm the capability of LLMs to evaluate other LLMs, with verification abilities improving faster than generation quality[29]. GPT-4’s evaluations exceeds 80% consistency with human reviewers[30] and exceeds 85% alignment in pairwise comparisons[31], reaching nearly 100% agreement when performance differences are significant[31]. Self-verification benefits LLM performance[32], though biases issues[33], such as position bias[34], length bias[30], and self-bias[35], persist but can be reduced through proper design[36].
Given characteristic writing patterns of LLMs and potential human evaluation bias, this study employs LLM-based evaluation exclusively. Existing LLM-based LLM evaluation methods encompass scoring, comparison, selection judgment, and comprehensive description. This study introduces a dual-baseline review quality evaluation framework, to minimize potential LLM evaluation bias and quantitatively compare LLM-generated reviews with peer-reviewed expert-written content, validating reliability through statistical analysis. In our method, we segmented 14 published Q1 reviews into 89 fragments based on semantic content. Using extracted topics from these fragments and the literature cited in the original text, we generated comparative reviews using Qwen2-7b-Instruct, Qwen2-72b-Instruct, and Claude3.5Sonnet. This methodology enabled direct comparison between human experts and LLMs in writing reviews with identical literature background, establishing a rigorous benchmark for LLMs given their limitations in accessing both human accumulated domain expertise and pre-1970 undigitized literature.
In the evaluation process, we compared the performance of two models of different scales from the open-source Qwen2 series (Qwen2-7b-Instruct and Qwen2-72b-Instruct) and the closed-source model Claude3.5Sonnet (Fig. 1). Evaluation employed both self-scoring and uniform scoring by Qwen2-72b-Instruct. Intraclass Correlation Coefficient (ICC) tests and Transitive Consistency Ratio (TCR) analyses confirmed high reliability for Claude3.5Sonnet (Fig. 1(a,b,e,f)) and Qwen2-72b-Instruct (Fig. 1(c,d,e,f)), meeting human evaluation standards, while Qwen2-7b-Instruct’s results necessitated substitution with Qwen2-72b-Instruct’s scoring due to insufficient reliability. This phenomenon may stem from differences in capabilities between LLMs of different scales, specifically manifested in three dimensions: world knowledge, language understanding, and logical reasoning. We believe that small models differ significantly from large models in logical reasoning ability, while world knowledge has been supplemented through context-provided literature, and language ability differences are relatively small. For details on Qwen2-7b-Instruct’s evaluation results, see SI.
Figure 1: Reliability verification results of the dual-baseline review quality assess-
ment framework. Scheme evaluation result distribution heat map, red to green showing 0% to +100% range, higher values indicate better performance: a, Intraclass correlation coefficient test for Claude3.5 Sonnet model results; b, Transitive consistency ratio analyses for Claude3.5 Sonnet model results; c, Intraclass correlation coefficient test for Qwen2-72b-Instruct model results; d, Transitive consistency ratio analyses for Qwen2-72b-Instruct model results. Scheme evaluation result distribution histograms: e, Intraclass correlation coefficient test results for Claude3.5 Sonnet model (green) and Qwen2-72b-Instruct model (red); f, Transitive consistency ratio analyses results for Claude3.5 Sonnet model (green) and Qwen2-72b-Instruct model (red). More data charts are available in the SI.
Model capability significantly impacts generation quality, while our method ensures a basic lower limit of generation quality. In repeated generation tests (9 times per model), average scores, taken as comprehensive performance of models, showed that Qwen2-7b-Instruct reached 43.94% of manual scoring, Qwen2-72b-Instruct reached 64.81% (Fig. 2(d,f)), and Claude3.5Sonnet exceeding manual scores by 23.63% (Fig. 2(c,f)), all significantly higher than the baseline level of direct generation. Optimal performance analysis, which taking the paragraph with the highest total score among those generated by each model as the optimal paragraph, revealed best-paragraph scores of 89.07% for Qwen2-7b-Instruct, 92.64% for Qwen2-72b-Instruct (Fig. 2(b,f)), and 130.79% for Claude3.5Sonnet relative to manual scores (Fig. 2(a,f)). These results indicate multiple generations and selecting the best paragraph can improve performance of smaller models, while larger models maintain consistent quality. For optimal results, we recommend using larger models when possible; otherwise, multiple generations can enhance final performance in hardware-constrained scenarios. For details on Qwen2-7b-Instruct’s evaluation results, see SI.
Figure 2: Quality assessment results of automatically generated reviews. Heat map of the percentage difference in scores of review paragraphs generated by this method relative to human scores, red to green showing -100% to +100% range, higher values indicate better performance, values truncated to ±100% range, values exceeding are recorded as -100% and +100%: a, Highest scoring paragraph of Claude3.5 Sonnet model; b, Highest scoring paragraph of Qwen2-72b-Instruct model; c, Average paragraph score of Claude3.5 Sonnet model; d, Average paragraph score of Qwen2-72b-Instruct model. e, Histogram of percentage differences in scores relative to human scores for highest scoring paragraphs, average paragraph scores, and directly generated paragraph scores without going through this method for Claude3.5 Sonnet model and Qwen2-72b-Instruct model, colors ranging from dark to light representing Claude3.5
Sonnet model and Qwen2-72b-Instruct model and Qwen2-72b-Instruct model, green, red, and blue representing highest scoring paragraphs, average paragraph scores, and directly generated result scores respectively. f, Bar chart of scores for highest scoring paragraphs of Claude3.5 Sonnet model and Qwen2-72b-Instruct model in each scoring sub-item, green and red representing Claude3.5 Sonnet model and Qwen2-72b-Instruct model respectively. More data charts are available in the SI.
The near-human scores of optimal paragraphs might reflect quality-related bias[37] because their closely approximate human-level quality could make LLMs’ potential evaluation bias more prominent, while the distinctly inferior quality of direct generation remains easily distinguishable. Analysis of optimal paragraphs reveals high correlations between Qwen2-7B and Qwen2-72B across evaluation sub-items (0.926 for highest scores, 0.939 for scores relative to human benchmarks), reflecting the scaling law of LLM and suggesting potential for further improvements with model advancement (see SI for details). Our research validates the effectiveness of LLM-based automated review generation, with quality approaching or exceeding manual reviews. The effectiveness of method stems from general language processing and context learning capabilities of LLMs, rather than requiring specific domain expertise, suggests broad disciplinary applicability. While inheriting common biases and requiring readers’ professional judgment, this method serves as a supportive rather than a replacement tool for human innovation, with open-source models showing comparable capabilities to closed-source alternatives. The approach demonstrates broad cross-disciplinary potential, promising to become an important tool for promoting academic innovation and knowledge dissemination, while the dual-baseline framework offers potential methodology for evaluating LLM agent workflows where manual data acquisition is costly.
@@ 2.4 Data mining and visual analysis
Catalysts play a vital role in chemical process optimization[38], with data mining enabling accelerated design through pattern recognition[39]. Analyzing 839 PDH catalyst papers from a total of 1041 articles filtered by abstracts and titles in Q1-Q3 chemistry journals (1980-2024), our data mining module revealed comprehensive insights into catalysts’ composition, structure, and performance.
For instance, publication analysis showed surging alloy research since 1995 and single-atom catalyst studies post-2015 (Fig. 3a), driven by advancements in structural composition (Fig. 3b). Performance analysis identified optimal promoter elements (Zn, Sn, La) (Fig. 3c) and support materials (alumina, zeolites) (Fig. 3d), while combination studies revealed superior performance in multi-metal systems, particularly Pt-based catalysts with Sn, Zn, In promoters (Fig. 3e). Moreover, impregnation-prepared nanometallic catalysts demonstrated high conversion rates and selectivity, contrasting with single-atom alloys’ high selectivity but lower conversion rates (Fig. 3f).
This comprehensive analysis reveals variable interactions and guides catalyst optimization, recommending Pt-based systems for selectivity and metal oxides for conversion rates, while highlighting the promise of single-atom and nanostructured catalysts. These findings not only establish performance benchmarks in PDH catalysis, but also demonstrate how our LLM-based methodology enables real-time scientific insight extraction, facilitating industrial-oriented catalyst design optimization.
Figure 3: Example of visual analysis results. Line charts for annual publication numbers: a, different catalyst types; b, Performance enhancement sources. Radar charts for peak performance of single factors, with selectivity (black) and stability (purple) scales: c, Promoter elements; d, Support materials. Bubble charts for dual-variable correlations, show selectivity (color depth), conversion rate (bubble size), and stability (bubble edge thickness), aiming for high selectivity, conversion rate, and stability. Data includes only those with selectivity $\geq$85%, conversion rate $\geq$45%, stability $\geq$1h: e, Active site element-composition element; f, Alloy structure type-preparation method. Complete data charts are available in the SI.